A rectangular corral needs to be built using 200feet of fencing. The formula for the corral area is A=100w-w^2 where A is the area and w is the width of the rectangle. Find the maximum possible area of the corral.
HELP! I need the answer please and thank you (:

Answer by Cromlix(4381)

A = 100w - w^2
Differentiate
A' = 100 - 2w
A' = 0
100 - 2w = 0
- 2w = - 100
2w = 100
w = 50
Using a nature table.
50 feet is the maximum width
Given 200 feet of fencing
Length = 50 feet
Width = 50ft
Area = length x width
50 x 50
2500 ft^2
Hope this helps.
:-)